Overhead Application, Overhead Variances, Journal Entries

Plimpton Company produces countertop ovens. Plimpton uses a standard costing system. The standard costing system relies on direct labor hours to assign overhead costs to production. The direct labor standard indicates that two direct labor hours should be used for every oven produced. The normal production volume is 100,000 units. The budgeted overhead for the coming year is as follows:

Fixed overhead $770,000
Variable overhead 446,000*
*At normal volume.

Plimpton applies overhead on the basis of direct labor hours.

During the year, Plimpton produced 97,000 units, worked 196,000 direct labor hours, and incurred actual fixed overhead costs of $780,200 and actual variable overhead costs of $437,540.

Required:

1. Calculate the standard fixed overhead rate and the standard variable overhead rate. Round your answers to the nearest cent. Use rounded answers in the subsequent computations.

Standard fixed overhead rate $ per direct labor hour
Standard variable overhead rate $ per direct labor hour

2. Compute the applied fixed overhead and the applied variable overhead. Use the application rates from part (1) in your calculations.

Fixed $
Variable $

What is the total fixed overhead variance?
$   

What is the total variable overhead variance?
$   

3. Break down the total fixed overhead variance into a spending variance and a volume variance.

Spending Variance $
Volume Variance $

4. Compute the variable overhead spending and efficiency variances.

Spending Variance $
Efficiency Variance $

5. Now assume that Plimpton’s cost accounting system reveals only the total actual overhead. In this case, a three-variance analysis can be performed. Using the relationships between a three- and four-variance analysis, indicate the values for the three overhead variances.

Volume variance $
Variable overhead efficiency variance $
Spending variance $

6. Prepare journal entries (1) to apply overhead to production, (2) to record the actual overhead costs incurred, (3) to record the variable and fixed overhead variances, and (4) to close the variance accounts at the end of the year. Assume variances are closed to Cost of Goods Sold. If an amount box does not require an entry, leave it blank or enter "0".
Answer #1

Answer:

Normal production units= 100,000
Standard labour hours= 2*100,000= 200,000 hrs.
1) Standard fixed overhead rate = std. fixed overhead/ std. labour hrs. = 770,000/200,000=3.85 per direct labour hrs.
Standard variable overhead rate= std. variable overhead/ std. labour hrs.= 446,000/200,000=2.23 per direct labour hrs.
2) Applied fixed overhead = std. fixed overhead rate* actual hours worked= 3.85*196,000=754,600
Applies variable overhead= std. var. overhead rate* actual hrs. woeked=2.23*196,000=437,080
Total fixed overhead variance = Actual fixed overhead - ( std. hrs.*actual production*fixed overhead absorption rate) = 780,200-(2*97,000*3.85)= 33,300 (unfavorable)
total variable overhead variance = Actual variable overhead - (std. hrs*actual production* variable overhead absorption rate) = 437,540-(2*97,000*2.23) = 4,920 (unfavorable)
3) Fixed overhead spending variance = Actual fixed overhead - ( std. fixed overhead rate*actual hrs.) = 780,200-(3.85*196,000)= 25,600 (unfavorable)
Fixed overhead volume variance= ( budgeted hrs. -actual hrs.) std. rate= (2*97,000-196,000)*3.85= 7,700 (unfavorable)
4) Variable overhead spending variance= Actual variable overhead -( std. variable overhead rate*actual hrs.) =
437,540-( 2.23*196,000)= 460 (unfavorable)
Variable overhead efficiency variance= (actual hrs. - std. hrs.)* std. overhead rate=
(196,000-2*97,000)*2.23= 4,460 ( unfavorable)